OpenStudy (anonymous):

Find the solution using Gaussin Elimination method: x-2y+z=6 y+2z=-7 x+y+3z=-7

OpenStudy (anonymous):

step 1) add all three equations step 2) subtract the second equation from the third equation step 3) post what you got here.

OpenStudy (anonymous):

okay so from adding up the three equations, i got 2x+2y+6z=-8. is that what i am subtracting from the second equation from? or is it the original 3rd equation i am using

OpenStudy (anonymous):

so you added the three equations is it? what is -2y + y +y ?

OpenStudy (anonymous):

okay so from adding up the three equations, i got 2x+2y+6z=-8. is that what i am subtracting from the second equation from? or is it the original 3rd equation i am using

OpenStudy (anonymous):

0

OpenStudy (anonymous):

then how did you get 2y when you added the three equations?

OpenStudy (anonymous):

okay so from adding up the three equations, i got 2x+2y+6z=-8. is that what i am subtracting from the second equation from? or is it the original 3rd equation i am using

OpenStudy (anonymous):

oh i read the last equation wrong at first

OpenStudy (anonymous):

okay. follow the steps and do the calculations correctly and post what you get here. also, in the second step, subract the original 2nd equation from the original third equation.